OpenStudy (anonymous):

solve the equation 8/x=5/4x+1

OpenStudy (anonymous):



OpenStudy (anonymous):

by just simplifying it

OpenStudy (anonymous):

with heros problem you'll distribute right?

OpenStudy (anonymous):

sorry I meant 8/x=5/(4X)+1

hero (hero):

Yes, you did. Now you know to include parentheses when posting

hero (hero):

\[\frac{8}{x} = \frac{5}{4x} + 1\] Perhaps this is what you meant.

hero (hero):

Multiply the fraction on the left by 4/4

OpenStudy (anonymous):

so now its

hero (hero):

@amandasoto144 ...NO